Hello,
Please be kind enough to help me with this question.
(w^a)^b * 16(w^b)^a
Thank You

Answer by user_dude2008(1862)

(w^a)^b * 16(w^b)^a = w^(ab)*16w^(ba) = w^(ab)*16w^(ab)=16w^(ab+ab)=16w^(2ab)


Answer: (w^a)^b * 16(w^b)^a = 16w^(2ab)